OpenStudy (anonymous):

Over a 5-year period, a company reported annual profits of $8 million, $3 million, $2 million, and $9 million. In the fifth year, it reported a loss of $7 million. What was the mean annual profit? $ _____ million

OpenStudy (anonymous):

Does your book give you the definition of mean? If not, it is the sum of all the the data points divided by the number of data points. In this case it the sum of the profits, divided by the number of years.

OpenStudy (anonymous):

8+3+2+9+7=28/5=5.6

OpenStudy (anonymous):

very close. Take a look at the fifth year.

OpenStudy (anonymous):

the annual profit

OpenStudy (anonymous):

The fifth year is a loss. You did everything right in your 2nd post except for one detail. How you interpret a "loss" mathematically?

OpenStudy (anonymous):

oh so its 4.2

OpenStudy (anonymous):

@undeadknight26

undeadknight26 (undeadknight26):

well u just add the numbers up and divide them by the amount of numbers there are.

OpenStudy (anonymous):

ya and it would be 4.2 right

undeadknight26 (undeadknight26):

i got 5.8...

OpenStudy (anonymous):

$8 is a loss

undeadknight26 (undeadknight26):

oh wait so: 8 + 3 + 2 + 9 - 7?

OpenStudy (anonymous):

indeed.

OpenStudy (anonymous):

i got 5.8 also but $8 is a loss

OpenStudy (anonymous):

8+3+2+8-7 all over 5

OpenStudy (anonymous):

sorry, 8+2+3+9-7

OpenStudy (anonymous):

so its actually 2.8

OpenStudy (anonymous):

i typed 8 a 2nd time by mistake.

OpenStudy (anonymous):

its 3.

OpenStudy (anonymous):

i am so confused because i divided it and got a different answer

OpenStudy (anonymous):

8+2+3+9 = 22. subtract 7, is 15. 15/5 is 3.

undeadknight26 (undeadknight26):

well it depends on what number u divide it by...if u include the number seven u will get a different score...im honestly confused my self because i kinda forgot this stuff..

OpenStudy (anonymous):

The mean just means average. you just add up the number from each year, and divide by the number of years. \[\frac{ 8+2+3+9+(-7) }{ 5 }\]

undeadknight26 (undeadknight26):

sounds straight.